# Auditing Human Decision-Making in High-Stakes Environments via Prescriptive AI: A Stress-Test on Real-Time Tactical Management

### v5

@@ 4.1 Baseline Limitations and Methodological Motivation
A central methodological challenge in this domain is quantifying player performance. The PlayeRank framework, proposed by Pappalardo2019, defines a multidimensional, role-aware metric that combines individual technical actions with influence on the passing network, validated against professional scout assessments. The authors explicitly aim to reduce biases related to playing time, normalizing performance vectors by the number of actions to “avoid biases due to play time.”
@@
Recent work supports this perspective. SchmidtLilloBustos2024 demonstrate that player influence fluctuates meaningfully over a match due to tactical reconfiguration, opponent pressure, and physical factors, advocating for temporal windowing to capture such dynamics. Following this insight, the present work also adopts 5-minute temporal slices, but advances the methodology by adding a role-aware percentile cumulative mean across slices. This modification preserves interpretability while enabling direct comparability between players with different playing times, eliminating residual exposure bias and allowing the performance curve to reflect true temporal variability.
@@
A second methodological challenge concerns the interpretation of imprecise or uncertain information in decision-making contexts. The literature highlights fuzzy logic as an effective tool for this purpose. For instance, Huarachi-Macuri et al. LACCEI2023 demonstrate that fuzzy systems are suitable for transforming quantitative descriptors into interpretable linguistic assessments (e.g., Stamina, Agility). Within tactical modeling, Marliere2017 uses fuzzy control systems to arbitrate between tactical states under uncertainty. These contributions motivate the role of fuzzy logic in the present system: integrating performance indicators and contextual factors into a Substitution Priority score in a manner compatible with the inherent uncertainty of in-game decision making.
@@
The decision to make substitutions in soccer remains an intrinsically human domain, hardly replicable by machine learning (ML) algorithms. While ML can analyze historical performance data, a coach’s choice involves a complex assessment of contextual, tactical, and psychological factors that are not easily quantifiable—such as team morale, perceived fatigue, reading the opponent’s strategy, and game momentum. Attempts to predict substitutions using ML, while informative, reveal a clear ceiling in their ability to capture this complexity. A recent study by mohandas, for example, analyzed a large dataset of 51,738 substitutions using multiple algorithms, including Random Forest, SVM, and Decision Trees. Even with this wealth of data, the best-performing model (Random Forest) achieved a maximum accuracy of just over 70%. This gap of approximately 30% is significant, suggesting that the final decision is not purely predictive but rather adaptive, relying on subjective human judgment that historical data alone cannot capture. Consequently, models attempting to predict the optimal substitution are conceptually limited. Alternative approaches, such as fuzzy logic-based decision support systems, appear more suitable, as they are designed not to replace but to support the coach’s judgment, managing the inherent uncertainty and qualitative factors of the match.
@@
Finally, the choice of an intrinsically interpretable methodology is not merely a design preference, but a methodological necessity in high-stakes decision-making contexts. As explicitly argued by Rudin2019, post-hoc explanations of black-box models are fundamentally inadequate when decisions have significant consequences, as they obscure the true decision logic and may provide misleading justifications. In domains such as elite soccer, where substitution decisions carry substantial sporting and financial impact, decision support systems must expose their reasoning process in a transparent and auditable manner. This perspective reinforces the limitations of purely predictive machine learning approaches discussed above and provides a principled justification for abandoning black-box optimization in favor of symbolic reasoning. Accordingly, the present work adopts fuzzy logic as the core inference mechanism—rather than as an explanatory add-on—ensuring that substitution recommendations are interpretable, accountable, and directly aligned with domain knowledge and human tactical judgment (Rudin2019).
@@
This perspective finds foundational support in the framework of logical dynamics proposed by vanBenthem2011, which conceptualizes rational agency as a continuous process of information update rather than static optimization. In this context, player substitutions serve as a particularly clear instance of epistemic actions (vanBenthem2007): they do not merely alter the physical parameters of the match, but fundamentally restructure the space of plausible future game states under severe informational constraints. At the moment of decision, the coach operates at the boundary between “hard facts” (e.g., scoreline, remaining time, substitution limits) and “soft information” (e.g., perceived fatigue, tactical momentum). Once executed, a substitution irreversibly fixes certain constraints while reweighting the plausibilities regarding future play. Consequently, decision quality cannot be assessed solely by ex-post outcomes, but by whether the action constituted a rational epistemic update given the information available at time $t$. This framing aligns the proposed system with a dynamic view of rationality—where correctness is understood as the capacity for informed correction under uncertainty—establishing its role not as a predictor of future success, but as an auditor of the epistemic transition induced by the substitution.
@@ 4.2.1 Problem Formulation
Tactical decision-making in elite soccer operates under high uncertainty and significant financial stakes GE_PremierLeague; ESPN_Mundial. Despite the proliferation of granular data, the specific process of substitution decisions remains predominantly intuitive or reliant on descriptive statistics that fail to capture real-time performance decay. This creates two distinct problems: the Metric Exposure Bias and the Predictive Ceiling.
The Metric Exposure Bias: Existing frameworks for player evaluation, such as the original PlayeRank Pappalardo2019, typically utilize cumulative sum formulations. While effective for post-match rankings, these metrics introduce a temporal bias during live games: a player’s score monotonically increases with playing time, regardless of their minute-by-minute efficiency. This mathematical structure masks declining performance, as a fatigued player performing poorly in the $80^{th}$ minute may still have a higher total score than a high-impact substitute, rendering standard metrics insufficient for real-time substitution decisions.
The Predictive Ceiling of Machine Learning: Current computational approaches to substitutions rely heavily on Supervised Machine Learning (SML) models trained on historical data. As noted by mohandas, these models achieve a prediction accuracy plateau of approximately 70%. This ”ceiling” exists because SML models are designed to mimic human behavior, thereby learning and replicating the cognitive biases of coaches (e.g., status quo bias or delaying defensive changes). Consequently, these models validate historical decisions rather than optimizing future outcomes, failing to identify necessary substitutions that deviate from conservative human norms.
@@ 4.3 Dataset
The database selected for this project is the ”Soccer match event dataset”, a detailed public repository of soccer match events (Pappalardo2020). The choice of this dataset is based on its high granularity. The dataset details player-level actions and aggregated performance metrics, allowing for the modeling of in-game situations, which is an essential requirement for our system. We used the complete dataset available on Kaggle (Pappalardo2020), which comprises 27 interrelated CSV tables. They cover seven main competitions and can be grouped into five logical categories:
@@
The choice of playerankScore as our main ”Performance” input is a central methodological decision. As proposed by Pappalardo2019, the PlayeRank framework was developed to solve the absence of a consolidated and universally accepted metric for evaluating player performance. The playerankScore is a metric derived from millions of game events that, according to the authors, surpasses other metrics when compared with assessments from professional scouts. Therefore, instead of trying to model performance from raw events, we adopted the playerankScore as an already validated and academically robust representation of a player’s performance in a match.
@@ 4.4 Data Integration and Pre-processing
The objective of this pipeline is to transform the raw multi-source event data (events, matches, players) into a single structured and temporal dataset tailored for the Fuzzy Logic Decision System. Unlike the original PlayeRank framework Pappalardo2019, which aggregates player performance per match, our system requires a fine-grained temporal perspective of the player’s performance evolution within a match. To achieve this, we implemented a three-stage processing pipeline that generates a dense temporal dataset, where each record represents the state of a player in a 5-minute (300-second) interval of play.
@@ 4.4.1 Stage 1: Temporal Performance Metric Generation
where $\alpha_{\mathrm{net}}=0.2$ controls the relative contribution of the network-based component. This configuration was selected to balance structural influence and technical efficiency, aligning with the sensitivity ranges reported in Pappalardo2019.
@@ 4.4.3 Stage 3: Final Cleaning and Role-Aware Normalization
  • Recalculation of Cumulative Means: The variable playerank_acumulativo_media_raw is recalculated using a robust expanding().mean() to ensure that averages reflect the final, harmonized temporal trace.
@@ 4.6.4 Inference and Final Calculation
Where $\alpha=0.25$ in the current version. This ensures that the fuzzy logic acts as a ”tuner” that intensifies or attenuates the urgency based on tactical context (e.g., a yellow card or a goal scored), rather than overriding the player’s core performance metric entirely. This methodology allows the expert’s knowledge ESPN_s_d; Terra_s_d; CNN_s_d to be faithfully captured by the system, adding significant weight to the player’s performance on the field.
4.7 A Normative Taxonomy of Action-Oriented AI Systems: Formalizing Prescriptive AI
This section establishes a coherent conceptual hierarchy for action-oriented decision systems. We first characterize the broad and heterogeneous superset of action-oriented systems. We then situate Prescriptive Analytics as a structured decision-oriented framework within that space. Finally, we formally delineate Prescriptive Artificial Intelligence (Prescriptive AI) as a normative, high-stakes subset of Prescriptive Analytics, defined not by algorithmic choices but by its epistemological role in human decision-making.
4.7.1 Action-Oriented Decision Systems: A Heterogeneous Superset
  • Optimization and mathematical programming systems;
  • Simulation-based and what-if analysis tools;
  • Rule-based and expert systems;
  • Reinforcement Learning (RL) agents;
  • Recommendation and next-best-action systems;
4.7.2 Prescriptive Analytics as a Decision-Oriented Subset
Within this heterogeneous landscape, Prescriptive Analytics provides a more structured and widely recognized framework. Originating in business intelligence and operations research, prescriptive analytics focuses on determining what actions should be taken to achieve specified objectives, typically through optimization, simulation, and rule-based reasoning Davenport2007; Lepenioti2020; Bertsimas2020.
Prescriptive Analytics narrows the scope of action-oriented systems by explicitly addressing decision recommendation. However, it remains largely agnostic with respect to how uncertainty, human judgment, or accountability should be handled. It encompasses both automated and decision-support systems and does not inherently distinguish between optimizing outcomes and supporting human deliberation.
Accordingly, Prescriptive Analytics is best understood as a decision-oriented subset of action-oriented systems: analytically grounded, but normatively under-specified.
Action-Oriented AI Systems (The Heterogeneous Superset) Reinforcement Learning (Automation-Oriented) Recommender Systems Predictive Workflows Prescriptive Analytics (Decision-Oriented) Black-Box Optimization Simulation-Based Planning Prescriptive AIDecision AuditingIntrinsic InterpretabilityRisk-Aware Reasoning
Figure 5: A normative taxonomy of action-oriented AI systems. The diagram illustrates the hierarchical relationship where Action-Oriented Systems form the superset. Prescriptive Analytics sits within as a decision-focused subset, while Prescriptive AI serves as the normative core, defined by auditability and interpretability requirements.
4.7.3 Formal Definition of Prescriptive AI
@@
Prescriptive AI systems are not characterized by the techniques they employ, but by the role they play in the decision process. In particular, they are distinguished by the following defining properties:
@@
  • Intrinsic interpretability as a structural requirement, not a post-hoc feature Rudin2019;

### v6

@@ 5.1 Baseline Limitations and Methodological Motivation
A central methodological challenge in this domain is quantifying player performance. The PlayeRank framework, proposed by Pappalardo et al. [30], defines a multidimensional, role-aware metric that combines individual technical actions with influence on the passing network, validated against professional scout assessments. The authors explicitly aim to reduce biases related to playing time, normalizing performance vectors by the number of actions to “avoid biases due to play time.”
@@
Recent work supports this perspective. Schmidt et al. [42] demonstrate that player influence fluctuates meaningfully over a match due to tactical reconfiguration, opponent pressure, and physical factors, advocating for temporal windowing to capture such dynamics. Following this insight, the present work also adopts 5-minute temporal slices, but advances the methodology by adding a role-aware percentile cumulative mean across slices. This modification preserves interpretability while enabling direct comparability between players with different playing times, eliminating residual exposure bias and allowing the performance curve to reflect true temporal variability.
@@
A second methodological challenge concerns the interpretation of imprecise or uncertain information in decision-making contexts. The literature highlights fuzzy logic as an effective tool for this purpose. For instance, Huarachi-Macuri et al. LACCEI [20] demonstrate that fuzzy systems are suitable for transforming quantitative descriptors into interpretable linguistic assessments (e.g., Stamina, Agility). Within tactical modeling, Marliere [23] uses fuzzy control systems to arbitrate between tactical states under uncertainty. These contributions motivate the role of fuzzy logic in the present system: integrating performance indicators and contextual factors into a Substitution Priority score in a manner compatible with the inherent uncertainty of in-game decision making.
@@
The decision to make substitutions in soccer remains an intrinsically human domain, hardly replicable by machine learning (ML) algorithms. While ML can analyze historical performance data, a coach’s choice involves a complex assessment of contextual, tactical, and psychological factors that are not easily quantifiable—such as team morale, perceived fatigue, reading the opponent’s strategy, and game momentum. Attempts to predict substitutions using ML, while informative, reveal a clear ceiling in their ability to capture this complexity. A recent study by Mohandas et al. [26], for example, analyzed a large dataset of 51,738 substitutions using multiple algorithms, including Random Forest, SVM, and Decision Trees. Even with this wealth of data, the best-performing model (Random Forest) achieved a maximum accuracy of just over 70%. This gap of approximately 30% is significant, suggesting that the final decision is not purely predictive but rather adaptive, relying on subjective human judgment that historical data alone cannot capture. Consequently, models attempting to predict the optimal substitution are conceptually limited. Alternative approaches, such as fuzzy logic-based decision support systems, appear more suitable, as they are designed not to replace but to support the coach’s judgment, managing the inherent uncertainty and qualitative factors of the match.
@@
Finally, the choice of an intrinsically interpretable methodology is not merely a design preference, but a methodological necessity in high-stakes decision-making contexts. As explicitly argued by Rudin [39], post-hoc explanations of black-box models are fundamentally inadequate when decisions have significant consequences, as they obscure the true decision logic and may provide misleading justifications. In domains such as elite soccer, where substitution decisions carry substantial sporting and financial impact, decision support systems must expose their reasoning process in a transparent and auditable manner. This perspective reinforces the limitations of purely predictive machine learning approaches discussed above and provides a principled justification for abandoning black-box optimization in favor of symbolic reasoning. Accordingly, the present work adopts fuzzy logic as the core inference mechanism—rather than as an explanatory add-on—ensuring that substitution recommendations are interpretable, accountable, and directly aligned with domain knowledge and human tactical judgment [39].
@@
This perspective finds foundational support in the framework of logical dynamics proposed by van Benthem [48], which conceptualizes rational agency as a continuous process of information update rather than static optimization. In this context, player substitutions serve as a particularly clear instance of epistemic actions [47]: they do not merely alter the physical parameters of the match, but fundamentally restructure the space of plausible future game states under severe informational constraints. At the moment of decision, the coach operates at the boundary between “hard facts” (e.g., scoreline, remaining time, substitution limits) and “soft information” (e.g., perceived fatigue, tactical momentum). Once executed, a substitution irreversibly fixes certain constraints while reweighting the plausibilities regarding future play. Consequently, decision quality cannot be assessed solely by ex-post outcomes, but by whether the action constituted a rational epistemic update given the information available at time $t$. This framing aligns the proposed system with a dynamic view of rationality—where correctness is understood as the capacity for informed correction under uncertainty—establishing its role not as a predictor of future success, but as an auditor of the epistemic transition induced by the substitution.
@@ 5.2.1 Problem Formulation
Tactical decision-making in elite soccer operates under high uncertainty and significant financial stakes [14, 12]. Despite the proliferation of granular data, the specific process of substitution decisions remains predominantly intuitive or reliant on descriptive statistics that fail to capture real-time performance decay. This creates two distinct problems: the Metric Exposure Bias and the Predictive Ceiling.
The Metric Exposure Bias: Existing frameworks for player evaluation, such as the original PlayeRank [30], typically utilize cumulative sum formulations. While effective for post-match rankings, these metrics introduce a temporal bias during live games: a player’s score monotonically increases with playing time, regardless of their minute-by-minute efficiency. This mathematical structure masks declining performance, as a fatigued player performing poorly in the $80^{th}$ minute may still have a higher total score than a high-impact substitute, rendering standard metrics insufficient for real-time substitution decisions.
The Predictive Ceiling of Machine Learning: Current computational approaches to substitutions rely heavily on Supervised Machine Learning (SML) models trained on historical data. As noted by [26], these models achieve a prediction accuracy plateau of approximately 70%. This ”ceiling” exists because SML models are designed to mimic human behavior, thereby learning and replicating the cognitive biases of coaches (e.g., status quo bias or delaying defensive changes). Consequently, these models validate historical decisions rather than optimizing future outcomes, failing to identify necessary substitutions that deviate from conservative human norms.
@@ 5.3 Dataset
The database selected for this project is the ”Soccer match event dataset”, a detailed public repository of soccer match events [31]. The choice of this dataset is based on its high granularity. The dataset details player-level actions and aggregated performance metrics, allowing for the modeling of in-game situations, which is an essential requirement for our system. We used the complete dataset available on Kaggle [31], which comprises 27 interrelated CSV tables. They cover seven main competitions and can be grouped into five logical categories:
  • 1.
@@
The choice of playerankScore as our main ”Performance” input is a central methodological decision. As proposed by Pappalardo et al. [30], the PlayeRank framework was developed to solve the absence of a consolidated and universally accepted metric for evaluating player performance. The playerankScore is a metric derived from millions of game events that, according to the authors, surpasses other metrics when compared with assessments from professional scouts. Therefore, instead of trying to model performance from raw events, we adopted the playerankScore as an already validated and academically robust representation of a player’s performance in a match.
@@ 5.4 Data Integration and Pre-processing
The objective of this pipeline is to transform the raw multi-source event data (events, matches, players) into a single structured and temporal dataset tailored for the Fuzzy Logic Decision System. Unlike the original PlayeRank framework Pappalardo et al. [30], which aggregates player performance per match, our system requires a fine-grained temporal perspective of the player’s performance evolution within a match. To achieve this, we implemented a three-stage processing pipeline that generates a dense temporal dataset, where each record represents the state of a player in a 5-minute (300-second) interval of play.
@@ 5.4.1 Stage 1: Temporal Performance Metric Generation
where $\alpha_{\mathrm{net}}=0.2$ controls the relative contribution of the network-based component. This configuration was selected to balance structural influence and technical efficiency, aligning with the sensitivity ranges reported in Pappalardo et al. [30].
@@ 5.4.3 Stage 3: Final Cleaning and Role-Aware Normalization
  • 3.
    Recalculation of Cumulative Means: The variable
    playerank_acumulativo_media_raw is recalculated using a robust expanding().mean() to ensure that averages reflect the final, harmonized temporal trace.
  • 4.
@@ 5.6.4 Inference and Final Calculation
Where $\alpha=0.25$ in the current version. This ensures that the fuzzy logic acts as a ”tuner” that intensifies or attenuates the urgency based on tactical context (e.g., a yellow card or a goal scored), rather than overriding the player’s core performance metric entirely. This methodology allows the coach’s in field knowledge to be faithfully captured by the system, adding significant weight to the player’s performance on the field.
6 Results
6.1 Overcoming the Predictive Ceiling: Prescription vs. Prediction
Recent benchmarks in soccer analytics reveal a predictive ceiling for substitution modeling. Mohandas et al. [26] analyzed 51,738 substitutions using supervised learning algorithms such as Random Forest and SVM, achieving a maximum accuracy of approximately $70\%$. This plateau indicates that nearly 30% of tactical decisions are influenced by contextual nuances—such as fatigue thresholds or specific game plans—that remain opaque to black-box models.
Supervised learning approaches in this domain suffer from a fundamental normative limitation: they are trained to mimic human experts. If historical data encode risk aversion, delayed reactions, or other systematic biases, the resulting models replicate them rather than correct them. As summarized in Table 6, our proposed framework shifts the objective from minimizing prediction error to maximizing tactical utility, acting as an auditor rather than an imitator.
Theoretical considerations further reinforce this distinction. As established in Theorem 2, constructing or benchmarking against an additional predictive substitution model does not provide a meaningful comparison. Predictive systems, by design, optimize for behavioral imitation or outcome likelihood. Consequently, they evaluate decisions with respect to variables that are epistemically unavailable at the moment of commitment. As formally shown in Section 3.1.2, such systems are inherently incapable of auditing decision quality without reintroducing outcome bias. Any predictive baseline—regardless of architectural sophistication or empirical performance—would therefore inherit the same normative limitation identified in prior work [e.g., 26]: replication of historical human biases rather than their diagnosis.
From this perspective, the relevant scientific question is not whether a stronger predictor could be engineered, but whether prediction is an appropriate objective for evaluating decisions under uncertainty. Our prescriptive framework answers this question in the negative by evaluating choices according to normative coherence with the epistemic state at time $t$. Requiring predictive parity as a validation standard would constitute a category error, conflating behavioral forecasting with normative decision auditing.
Dimension Supervised ML Baseline (e.g., Mohandas et al. [26]) Proposed Prescriptive System
6.2 The Mathematical Mechanism: Eliminating Exposure Bias
6.3 Stress-Testing Decision Latency: The Brazil vs. Belgium Case
6.3.1 Alignment and Early Detection
  • 1.
    Willian (45’): Priority $P_{final}=72.0$, ranked 4th. Rule R04 (Rapid Drop) flagged a sharp decline in technical actions in the final 20 minutes of the first half. Immediate substitution by the coach confirms offensive ineffectiveness, supported by media reports [11].
  • 2.
    Gabriel Jesus (58’): Priority $P_{final}=99.1$, 2nd highest. Rule R08 (Ineffective Forward) triggered by low cumulative score and no goals/assists. Post-match data: 12 touches in 58 minutes [38], validating tactical isolation.
  • 3.
    Paulinho (73’): Priority $P_{final}=93.1$. System detected ”Early Fatigue” and low participation. External analysis noted minimal offensive impact [13].
6.3.2 The “Fagner Paradox”: Counterfactual Auditing
Table 7: Model Audit vs. Human Decision (Brazil vs. Belgium)
Player Analysis Slice($t$) $P_{\text{final}}$ Rank Diagnosis Real Decision
Willian 40’-45’ 72.0 4th Rapid Drop Substituted out
Gabriel Jesus 55’-60’ 99.1 2nd Ineffective Forward Substituted out
Fagner 45’-85’ 100.0 1st Critical Risk Not Substituted
Paulinho 65’-73’ 93.1 3rd Early Fatigue Substituted out
6.3.3 The “Lukaku Paradox”: Latency and Event Masking
  • 1.
  • 2.
Table 8: Comparative Analysis of Low-Volume Strikers
Player Context FCS Priority Human Response
G. Jesus (BRA) 12 Touches, 0 G/A High (#2) Substituted (58’)
R. Lukaku (BEL) 14 Touches, 1 Assist Max (#1) Delayed Sub (87’)
6.4 Systemic Validation: Longitudinal Patterns
Technical Anchors vs. Chronic Inefficacy.
Temporal Gap and Anticipation.
Post-Entry Validation.
6.5 Boundary Conditions: The Limits of Performance Auditing
7 Discussion
@@
  • 1.
@@
  • 3.
    Intrinsic interpretability as a structural requirement, not a post-hoc feature [39];
  • 4.